OpenStudy (anonymous):

10. Which of the following could NOT be the lengths of the sides of a right triangle? (1 point) 9 ft, 12 ft, 15 ft 5 in., 10 in., 15 in. 4 cm, 7.5 cm, 8.5 cm 1.5 m, 2 m, 2.5 m

OpenStudy (istim):

Try using the Pythagorean Theorem. c^2=a^2+b^2

OpenStudy (anonymous):

i still need help

OpenStudy (p0sitr0n):

exact. if the equality is not true, then it isnt a right triangle

OpenStudy (pfenn1):

look at each set of numbers. do they make sense? try to picture what the resulting triangles would look like....I'll give you a hint......you can't have the lengths two sides add up to the length of the third

OpenStudy (anonymous):

is it 5 in., 10 in., 15 in.

OpenStudy (p0sitr0n):

100 + 25 =125

OpenStudy (anonymous):

time out im lost

OpenStudy (pfenn1):

use the Pythagorean Theorem as shown above. 10 squared is 100, 5 squared is 25 but 15 squared is 225, not the sum of 100 = 25.

OpenStudy (pfenn1):

*100 + 25